INTRACAVITY LASER ABSORPTION SPECTROSCOPY OF FLAMES. HCO RADICAL DETECTION IN METHANE/AIR FLAME

\underline{Sergey Cheskis}, School of Chemistry, Sackler Faculty of Exact Sciences, Tel Aviv University, Ramat Aviv, Tel Aviv 69978, Israel; Vladimir A. Lozovsky, Institute of Chemical Physics, Russian Academy of Sciences, Kosygina str. 4, 117977 Moscow, Russia.

Formyl radical, HCO, was monitored for the first time in an atmospheric pressure premixed hydrocarbon flame Intracavity Laser Absorption Spectroscopy (ICLAS) based on quasi-cw argon-ion pumped dye laser was used. HCO absolute concentration profile was also measured in 30 torr flat methane/air flame. The ability of ICLAS to the absolute concentration and temperature measurements in flames is discussed. The sensitivity of the detection is 1·10^{12} cm^{−3} and can be improved with isolation of the laser cavity from atmospheric water background.